### v1

@@ 5 Conclusion
In this work, we proposed and validated the efficacy of ECGFounder in adapting to a wide range of cardiovascular diagnostic applications, demonstrating its high performance and versatility across various downstream tasks as a foundational ECG model. By overcoming the limitations related to ECG data and labeling diversity, as well as training methods, the ECG foundation model confirms its potential to transform the standard of care in cardiology and to provide real-time, accurate cardiac assessments in diverse clinical settings.
@@
Appendix A Details of Dataset
In this section, the comparison and split methods of the dataset are introduced in detail.
@@ Appendix A Details of Dataset
Table LABEL:tb:data is a supplement to Figure 1.c in the main text. The table compares the size and label types of ECG datasets used in multiple existing works. It can be observed that HEEDB has far more data sizes and label types than other datasets.
Table 3: Comparison with ECG dataset of HEEDB and existing works.
# ECG # Patients # Classes # Lead
Stanford-ECG [11] 91,232 53,549 12 1
PTB-XL [38] 21,837 18,885 71 12
PhysioNet Challenge 2020 [26] 66,361 Unknown 27 12
CODE [31] 2,322,513 1,676,384 6 12
Tongji-ECG [42] 180,112 70,692 21 12
UCSF-ECG [16] 992,748 365,009 38 12
HEEDB (Ours) [19] 10,771,552 1,818,247 150 12
Table 4: Age, working year and education level of the cardiologists participating in the comparison with ECGFounder.
Member Working year Age Education level
Student Cardiologist 1 2 25 M.D. Candidate
Student Cardiologist 2 3 27 M.D. Candidate
Student Cardiologist 3 1 24 M.D. Candidate
Expert Cardiologist 1 25 50 Master
Expert Cardiologist 2 27 48 Master
Table 5: Details on fine-tuning dataset split.
Dataset Task type Train Valid Test
HEEDB classification 6,683,587 835,448 834,926
MIMIC-ECG-NTproBNP classification 47,084 5,884 5,893
MIMIC-ECG-CKD classification 213,926 26,749 26,741
MIMIC-ECG-LVEF classification 60,333 7,539 7,541
MIMIC-ECG-CHD classification 213,926 26,749 26,741
MIMIC-ECG-Age classification 640,027 80,015 80,004
MIMIC-ECG-Sex classification 640,027 80,015 80,004
MIMIC-ECG-NTproBNP regression 47,084 5,884 5,893
MIMIC-ECG-LVEF regression 60,333 7,539 7,541
MIMIC-ECG-Age regression 640,027 80,015 80,004
Deepbeat(PPG) classification 466,904 51,878 17,617
Appendix B Details of Labels
Table 6: Filtered 150 meaningful label list of HEEDB development set.
@@ Appendix B Details of Labels
Appendix C Details of Model
@@
Figure 4: The detailed architecture plot of the ECGFounder, indicating the dimensions of key components of the model.
Table 7: Hyperparameters for pre-training.
Hyperparameters Values
Batch size 1024
Learning rate 1e-3
Minimal learning rate 1e-6
Learning rate scheduler Cosine
Optimizer AdamW
Weight decay 0.1
Total epochs 20
Table 8: Hyperparameters for fine-tuning.
Hyperparameters Values
Batch size 256
Learning rate 1e-4
Minimal learning rate 1e-6
Learning rate scheduler ReduceLROnPlateau
Optimizer AdamW
Weight decay 0.1
Total epochs 30
Details of Results
Table 9: Performance of deep learning model and cardiologists on the internal test set
Model Cardiologists
Class Count AUC Sens Spec Sens Spec
ANTERIOR INFARCT 2 0.908 1.000 0.906 0.750 0.999
ATRIAL FIBRILLATION 59 0.996 1.000 0.972 0.798 0.981
ATRIAL FLUTTER 19 0.993 1.000 0.986 0.568 0.990
ATRIAL-PACED RHYTHM 3 0.991 1.000 0.988 0.833 0.995
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 4 0.893 0.959 0.898 0.563 0.977
INFERIOR INFARCT 1 0.998 1.000 0.998 1.000 0.997
LATERAL INFARCT 2 0.998 1.000 0.998 0.250 0.999
LEFT BUNDLE BRANCH BLOCK 11 1.000 1.000 0.998 0.600 0.932
NORMAL SINUS RHYTHM 364 0.969 0.934 0.926 0.916 0.862
PREMATURE ATRIAL COMPLEXES 19 0.997 1.000 0.963 0.579 0.989
PREMATURE VENTRICULAR COMPLEXES 40 0.981 0.975 0.889 0.850 0.996
RIGHT AXIS DEVIATION 12 0.996 1.000 0.992 0.594 0.979
RIGHT BUNDLE BRANCH BLOCK 50 0.984 0.940 0.969 0.775 0.970
SINUS BRADYCARDIA 41 0.995 1.000 0.989 0.717 0.994
SINUS RHYTHM 191 0.970 0.970 0.926 0.916 0.862
SINUS TACHYCARDIA 58 0.996 1.000 0.945 0.821 0.990
VENTRICULAR TACHYCARDIA 9 0.903 1.000 0.882 0.875 0.996
VENTRICULAR-PACED RHYTHM 20 0.988 0.950 0.940 0.725 0.997
WITH 1ST DEGREE AV BLOCK 15 0.864 0.667 0.831 0.633 0.962
WITH SINUS ARRHYTHMIA 6 0.976 1.000 0.945 0.542 0.979
@@
Figure 5: ROC curves on the internal test set of the deep learning model, compared with cardiologists
@@
Table 10: Performance of other ECG deep learning model and ECGFounder on external test set (single-lead ECG) PhysioNet Challenge-2017
Models ECGFounder Stanford[11]
@@
ATRIAL FIBRILLATION 0.957 0.924 0.979 0.756 0.840
Figure 6: AUROC of single-lead ECG downstream tasks.
Figure 7: Regression scatter of ECG downstream tasks.
Appendix D Data link
Appendix E Model weights link

### v2

@@ 5 Conclusion
In this work, we proposed and validated the efficacy of ECGFounder in adapting to a wide range of cardiovascular diagnostic applications, demonstrating its high performance and versatility across various downstream tasks as a foundational ECG model. By overcoming the limitations related to ECG data and labeling quality and diversity, as well as training methods, our ECG foundation model confirms its potential to transform the standard of care in cardiology and to provide real-time, accurate cardiac assessments in diverse clinical settings. We have provided open access to the model, as well as the code and data used to train it under an open access license. In this way, we invite the research community to continue build upon our model, and help advance the state-of-the-art.
Acknowledgements
This document is the result of the research project funded by the National Science Foundation and Emory University via an unrestricted gift.
Retrospective analysis of data for this project was conducted with waiver of informed consent under approved IRB protocols (BIDMC: 2022P000417; MGH: 2013P001024).
Funding: Dr. Westover was supported by grants from the NIH (RF1AG064312, RF1NS120947, R01AG073410, R01HL161253, R01NS126282, R01AG073598, R01NS131347, R01NS130119), and NSF (2014431). Dr. Shenda Hong was supported by the National Natural Science Foundation of China (62102008), Clinical Medicine Plus X - Young Scholars Project of Peking University, the Fundamental Research Funds for the Central Universities (PKU2024LCXQ030). For this work, Dr. Clifford was partially supported by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under NIH award number R01EB030362.
Disclosures: Dr. Westover is a co-founder, scientific advisor, and consultant to Beacon Biosignals and has a personal equity interest in the company. Dr. Clifford holds significant stock in AliveCor.
The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health, or the authors’ current and past employers and funding bodies.
We thank Dr. Xinxin Di, Dr. Kun Lu, Dr. Zhengkai Xue, Dr. Wenbo Dai, Dr. Jing Zhao, Dr. Hongqian Zhou for annotating the internal test set.
@@ Acknowledgements
S1. Q&A of the ECGFounder
What is the ECG foundation model?
The ECG foundation model is a large-scale, pre-trained neural network designed for tasks involving ECG (electrocardiogram) data. This model aims to capture and learn generalizable features from ECG signals, which can be fine-tuned or adapted for a wide range of downstream tasks like disease classification, arrhythmia detection, or patient-specific anomaly identification.
In our study, we propose the ECGFounder, which is being developed using the Harvard-Emory ECG Dataset. The model leverages this diverse and comprehensive dataset to create a robust feature extractor that can later be fine-tuned or used in transfer learning scenarios for specific diagnostic or research objectives. The ultimate goal is to have a powerful, generalized model that can perform various ECG-related tasks with minimal additional training on specific datasets.
What is the use of ECG foundation model?
We provide the medical AI community with a versatile ECG foundation model that serves as a comprehensive feature extractor and a highly adaptable base for transfer learning tasks. Compared to traditional models trained from scratch, this foundation model offers a generalized representation of ECG signals, enabling enhanced performance across diverse downstream tasks through fine-tuning. By leveraging large-scale pre-training on the Harvard-Emory ECG Dataset, the model captures robust and clinically relevant features, thus reducing the need for extensive labeled datasets for specific applications. This advancement holds significant potential to drive future innovations in AI-based ECG analysis, disease detection, and personalized healthcare solutions. And we offer an effective out-of-the-box solution and fine-tuning capabilities for the community to use it more easily.
How is the performance of the ECG foundation model?
The ECG foundation model demonstrates strong performance through several tasks, showcasing its high accuracy and generalization capabilities due to pre-training on large dataset like the Harvard-Emory ECG Dataset. It achieves superior results across various downstream tasks, including arrhythmia classification, disease diagnosis, and demographics detection, with fine-tuning further enhancing performance compared to models trained from scratch. The model effectively captures important ECG features, enabling precise differentiation of cardiac conditions with minimal domain-specific training. Additionally, its pre-trained architecture allows for efficient transfer learning, requiring fewer labeled samples and training epochs, thereby reducing time and computational costs. Its adaptability extends beyond disease classification to tasks like demographics and laboratory measurement detection, highlighting its versatility. Validation results consistently demonstrate that the foundation model outperforms baseline ECG models, achieving higher AUC, making it a robust and efficient tool for advancing AI applications in cardiology and beyond. The table S1 and S2 show the performance of ECG foundation model.
Table S1: Performance of classification tasks of the ECG foundation model.
Dataset Task type ECGFounder ECG-SimCLR ECG-ResNet
MIMIC-ECG-NTproBNP Laboratory measurement 0.8504 0.8133 0.8007
MIMIC-ECG-CKD Disease diagnosis 0.7459 0.7189 0.7093
MIMIC-ECG-LVEF Medical image metric 0.8674 0.8552 0.8543
MIMIC-ECG-CHD Disease diagnosis 0.7951 0.7631 0.7550
MIMIC-ECG-Age Demographics 0.8613 0.8394 0.8033
MIMIC-ECG-Sex Demographics 0.8983 0.8736 0.8412
Deepbeat(PPG) PPG AF detection 0.8203 0.7984 0.7950
Table S2: Performance of regression tasks of the ECG foundation model.
Dataset Task type ECGFounder ECG-SimCLR ECG-ResNet
MIMIC-ECG-NTproBNP Laboratory measurement 3.8468 4.8873 4.9893
MIMIC-ECG-LVEF Medical image metric 7.0117 7.0579 7.9514
MIMIC-ECG-Age Demographics 8.6501 9.0153 9.4004
How can researchers use the ECG Foundation Model?
The trained model and labeled data will be publicly released upon publication through the Brain Data Science Platform (bdsp.io), huggingface and PhysioNet.
S2. Details of Datasets
S2.1. Details of pre-training dataset
@@
Table LABEL:tb:data is a supplement to Figure 1.c in the main text. The table compares the size and label types of ECG datasets used in multiple existing works. It can be observed that Harvard-Emory ECG Database(HEEDB) has far more data sizes and label types than other datasets. And table 2 describes the label types and corresponding quantities of all HEEDB datasets.
Table S3: Comparison with ECG dataset of HEEDB and existing works.
ECGs Patients Classes Lead
Stanford-ECG 32 91,232 53,549 12 1
PTB-XL 18 21,837 18,885 71 12
PhysioNet Challenge 2020 37 66,361 Unknown 27 12
CODE 17 2,322,513 1,676,384 6 12
Tongji-ECG 6 180,112 70,692 21 12
UCSF-ECG 33 992,748 365,009 38 12
HEEDB (Ours) 15 10,771,552 1,818,247 150 12
Table S4: Filtered 150 meaningful label list of HEEDB development set.
@@
S2.2. Details of fine-tuning dataset
To evaluate the capability of transfer learning of ECGFounder, we developed 10 kinds of downstream datasets for fine-tuning.
For the MIMIC-ECG-NTproBNP dataset, we extracted data from the lab events section of MIMIC-IV-Hosp, obtaining NT-proBNP values matched with corresponding patients and admission IDs. According to clinically recognized laboratory standards, NT-proBNP values were categorized into two groups: normal and abnormal. This classification, along with the raw NT-proBNP values, was then matched to the corresponding ECGs of the patients for use in classification and regression tasks, respectively. Additionally, considering that abnormal NT-proBNP values can sometimes be excessively high, leading to non-convergence in the regression loss, we applied z-score normalization to these values. All normalized values were scaled to a range of 0 to 50.
For the MIMIC-ECG-CKD dataset, we extracted data from the diagnoses-ICD section of MIMIC-IV-Hosp, obtaining ICD codes matched with corresponding patients and admission IDs. Considering that the MIMIC database contains both ICD-9 and ICD-10 formats, we selected ICD-10 as the reference standard and converted all ICD codes to ICD-10. The ICD-10 codes for chronic kidney disease (CKD) include N182 (Chronic kidney disease, stage 2 [mild]), N183 (Chronic kidney disease, stage 3 [moderate]), N184 (Chronic kidney disease, stage 4 [severe]), N185 (Chronic kidney disease, stage 5), and N186 (End-stage renal disease). These codes were consolidated and treated collectively as indicators of CKD. If a patient has a positive diagnosis under any of these codes, they are considered to have CKD. This CKD classification information was then matched to the corresponding ECGs of the patients for use in the CKD classification task.
For the MIMIC-ECG-CHD dataset, we extracted ICD codes following the same procedure as described above. The ICD-10 codes for chronic heart disease (CHD) include all codes from I20 to I25. These codes were consolidated and treated collectively as indicators of CHD. If a patient has a positive diagnosis under any of these codes, they are considered to have CHD. This CHD classification information was then matched to the corresponding ECGs of the patients for use in the CHD classification task.
For the MIMIC-ECG-LVEF dataset, we extracted data from the discharge section of MIMIC-IV-Notes, obtaining echocardiogram notes and corresponding LVEF values matched with patients and admission IDs. According to clinically recognized echocardiogram diagnostic standards, an LVEF of 50% or higher is considered normal, while an LVEF below 50% is considered abnormal. This LVEF classification information, along with the raw LVEF values, was matched to the corresponding ECGs of the patients for use in both classification and regression tasks.
For the MIMIC-ECG-Age dataset, we extracted data from the patients section of MIMIC-IV-Hosp, obtaining age information matched with the corresponding patients. Ages were classified using a threshold of 65 years: individuals aged 65 or older were considered elderly, while those under 65 were not. This age classification, along with the raw age values, was matched to the corresponding ECGs of the patients for use in both classification and regression tasks.
For the MIMIC-ECG-Sex dataset, we extracted data from the patients section of MIMIC-IV-Hosp, obtaining sex information matched with the corresponding patients. This sex classification information was then matched to the corresponding ECGs of the patients for use in the sex classification task.
For the PPG Dataset, we utilized the DeepBeat-PPG dataset, which includes PPG data from 132 patients with atrial fibrillation (AF). This dataset provides publicly available validation and test sets. Here, we used the validation set as both the training and validation set, and the test set as a holdout test set. This approach ensures that patients do not overlap between the training and test sets, allowing for a fair performance comparison.
Table S5: Details on pre-training and fine-tuning dataset split.
Dataset Task type Train Valid Test
HEEDB(Pre-training) classification 6,683,587 835,448 834,926
MIMIC-ECG-NTproBNP classification 47,084 5,884 5,893
MIMIC-ECG-CKD classification 213,926 26,749 26,741
MIMIC-ECG-LVEF classification 60,333 7,539 7,541
MIMIC-ECG-CHD classification 213,926 26,749 26,741
MIMIC-ECG-Age classification 640,027 80,015 80,004
MIMIC-ECG-Sex classification 640,027 80,015 80,004
MIMIC-ECG-NTproBNP regression 47,084 5,884 5,893
MIMIC-ECG-LVEF regression 60,333 7,539 7,541
MIMIC-ECG-Age regression 640,027 80,015 80,004
Deepbeat(PPG) classification 466,904 51,878 17,617
S2.3. Details of cardiologists
Table S6 describes the personal information of 5 cardiologists participating in Cardiologist versus ECGFounder, including working year, gender, age and education level.
Table S6: Working year, gender, age and education level of the cardiologists participating in the comparison with ECGFounder.
Member Working year Gender Age Education level
Student Cardiologist 1 2 Male 25 M.D. Candidate
Student Cardiologist 2 3 Female 27 M.D. Candidate
Student Cardiologist 3 1 Male 24 M.D. Candidate
Expert Cardiologist 1 27 Female 50 Master
Expert Cardiologist 2 25 Male 48 Master
S3. Details of Model
S3.1. Details of model architecture
@@
Figure S1: The detailed architecture plot of the ECGFounder, indicating the dimensions of key components of the model.
In the input layer of ECGFounder, the model receives a multi-channel ECG signal represented by a size of 12×5000, indicating 12 channels and 5000 time steps. This input passes through a 1D convolutional layer with an input size of 12×5000, where 64 filters of size 64×64 are applied to produce a feature map with dimensions of 64×160. As the network progresses beyond the first layer, multiple convolutional blocks are introduced, with each block composed of convolution and pooling operations that gradually increase the network depth. Initially, these transformations reduce the feature map size from 64×160 to 32×80, eventually leading to an output size of 8×400, achieved through two repetitions of the convolutional block.
Subsequently, from the third to the fifth layer, the network continues to apply convolution and pooling, which results in further dimensional reduction while progressively extracting more intricate features. The third layer’s output remains at 400×400, with three repetitions of the convolution operation, while the fourth layer expands to produce a feature map of 1024×1024, repeated four times. Following these layers of convolution and feature extraction, the network transitions to a fully connected layer. Here, the extracted features from the preceding layers serve as input, and the fully connected operation generates either a single value or a categorical prediction based on the task at hand.
Notably, skip connections are included between layers to mitigate the vanishing gradient problem typically encountered in deep networks. These connections not only improve the training process but also enhance overall model performance by facilitating efficient information flow across the network.
S3.2. Details of training
The training process for ECGFounder involved distinct sets of hyperparameters for pre-training and fine-tuning stages, optimized to achieve robust model performance. During the pre-training phase, a large batch size of 1024 was employed alongside a relatively high learning rate of 1e-3, which was gradually reduced using a ReduceLROnPlateau scheduler to a minimum of 1e-5. The AdamW optimizer was utilized with a weight decay of 0.1 to enhance generalization, and training was conducted over a total of 20 epochs.
In the fine-tuning phase, the batch size was reduced to 256, and the learning rate was lowered to 1e-4 with a minimum threshold of 1e-6. The ReduceLROnPlateau scheduler was retained to adaptively manage the learning rate, while the AdamW optimizer and a weight decay of 0.1 continued to be employed. The total number of epochs was increased to 30, allowing the model to refine its performance on downstream tasks. This systematic approach to hyperparameter selection ensured the model’s adaptability and stability across various training stages.
S4. Details of Results
To further compare the performance of ECGFounder in clinical diagnosis with that of experts, we compared the performance of the model with that of 5 cardiologists. The results are shown in the figure S2 and tableLABEL:tab:internal.
Table S7: Performance of ECGFounder and cardiologists on the committee test set
ECGFounder Cardiologists
Class Count AUC Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 2 0.908 1.000 0.906 0.408 0.750 0.999 0.625
ATRIAL FIBRILLATION 59 0.996 1.000 0.972 0.866 0.798 0.981 0.796
ATRIAL FLUTTER 19 0.993 1.000 0.986 0.759 0.568 0.990 0.514
ATRIAL-PACED RHYTHM 3 0.991 1.000 0.988 0.500 0.833 0.995 0.657
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 4 0.893 1.000 0.698 0.506 0.563 0.977 0.299
INFERIOR INFARCT 1 0.998 1.000 0.998 0.667 1.000 0.997 0.725
LATERAL INFARCT 2 0.998 1.000 0.998 0.667 0.250 0.999 0.250
LEFT BUNDLE BRANCH BLOCK 11 1.000 1.000 0.998 0.909 0.600 0.932 0.371
NORMAL SINUS RHYTHM 364 0.969 0.934 0.926 0.952 0.916 0.862 0.930
PREMATURE ATRIAL COMPLEXES 19 0.997 1.000 0.963 0.679 0.579 0.989 0.613
PREMATURE VENTRICULAR COMPLEXES 40 0.981 0.975 0.889 0.600 0.850 0.996 0.898
RIGHT AXIS DEVIATION 12 0.996 1.000 0.992 0.800 0.594 0.979 0.489
RIGHT BUNDLE BRANCH BLOCK 50 0.984 0.940 0.969 0.847 0.775 0.970 0.762
SINUS BRADYCARDIA 41 0.995 1.000 0.989 0.938 0.717 0.994 0.791
SINUS RHYTHM 191 0.970 0.970 0.926 0.971 0.916 0.862 0.930
SINUS TACHYCARDIA 58 0.996 1.000 0.945 0.683 0.821 0.990 0.833
VENTRICULAR TACHYCARDIA 9 0.903 1.000 0.882 0.635 0.875 0.996 0.678
VENTRICULAR-PACED RHYTHM 20 0.988 0.950 0.940 0.559 0.725 0.997 0.756
WITH 1ST DEGREE AV BLOCK 15 0.864 0.667 0.831 0.287 0.633 0.962 0.560
WITH SINUS ARRHYTHMIA 6 0.976 1.000 0.945 0.308 0.542 0.979 0.319
@@
Figure S2: ROC curves on the internal test set of the deep learning model, compared with cardiologists
We evaluated single-lead atrial fibrillation detection on an external validation set. As shown in the tableS8, the performance of our model exceeds the previous SOTA method.
@@
Table S8: Performance of other ECG deep learning model and ECGFounder on external test set (single-lead ECG) PhysioNet Challenge-2017
Models ECGFounder Stanford32
@@
ATRIAL FIBRILLATION 0.957 0.924 0.979 0.856 0.840
Moreover, as shown in the figure S3 and S4, we evaluated several downstream tasks on single-lead ECGFounder thorough fine-tuning.
Figure S3: AUROC of single-lead ECG downstream tasks.
Figure S4: Regression scatter of single-lead ECG downstream tasks.
S5. Data link
S6. Model weights link